### v1

# On The Fragility of Benchmark Contamination Detection in Reasoning Models

Abstract
Leaderboards for large reasoning models (LRMs) have turned evaluation into a competition, incentivizing developers to optimize directly on benchmark suites. A shortcut to achieving higher rankings is to incorporate evaluation benchmarks into the training data, thereby yielding inflated performance, known as benchmark contamination. Despite that numerous contamination detection approaches have been proposed, surprisingly, our studies find that evading contamination detections for LRMs is alarmingly easy. We focus on the two scenarios where contamination may occur in practice: (I) when the base model evolves into LRM via supervised fine-tuning (SFT) and reinforcement learning (RL), we find that contamination during SFT can be originally identified by contamination detection methods. Yet, even a brief Group Relative Policy Optimization (GRPO) training can markedly conceal contamination signals that most detection methods rely on. Further empirical experiments and theoretical analysis indicate that Proximal Policy Optimization (PPO) style importance sampling and clipping objectives are the root cause of this detection concealment, indicating that a broad class of RL methods may inherently exhibit similar concealment capability; (II) when SFT contamination with CoT is applied to advanced LRMs as the final stage, most contamination detection methods perform near random guesses. Without exposure to non-members, contaminated LRMs would still have more confidence when responding to those unseen samples that share similar distributions to the training set, and thus, evade existing memorization-based detection methods. Together, our findings reveal the unique vulnerability of LRMs evaluations: Model developers could easily contaminate LRMs to achieve inflated leaderboards performance while leaving minimal traces of contamination, thereby strongly undermining the fairness of evaluation and threatening the integrity of public leaderboards. This underscores the urgent need for advanced contamination detection methods and trustworthy evaluation protocols tailored to LRMs. Our code is available at https://github.com/ASTRAL-Group/LRM_Conta_Detection_Arena.git.
1 Introduction
Competition among model developers has intensified as Large Language Models (LLMs) have demonstrated remarkable capabilities in various real-world tasks (Achiam et al., 2023; Wang et al., 2024). The leaderboards for performance are becoming a competitive arena for all state-of-the-art (SOTA) LLMs. However, inadvertently, benchmark samples may appear during LLMs’ pre-training due to vast amounts of web-scraped training data. In addition, in the pursuit of publicity, some model developers may even deliberately incorporate benchmark data into their training sets (Sun et al., 2025), resulting in inflated benchmark performance and leaderboard rankings. We refer to this as the benchmark contamination problem in LLMs (Xu et al., 2024; Balloccu et al., 2024).
Accordingly, various benchmark contamination detection methods have been proposed to determine whether specific benchmarks were used during training (Yeom et al., 2018; Mattern et al., 2023; Shi et al., 2023; Dong et al., 2024; Tu et al., 2024), based on the assumption that contamination in LLMs primarily involves memorizing the benchmark data (Wu et al., 2025). These methods rely on separability in some distributions between members (i.e., seen samples during contamination) and non-members (i.e., unseen samples). However, as LLMs have started to evolve into Large Reasoning Models (LRMs) (Guo et al., 2025; Jaech et al., 2024), benchmark contamination detection faces two key challenges: (1) LRMs rely on chain-of-thought (CoT) reasoning to reach final answers, but model developers would not release their training CoT data, and contamination detectors typically only have access to question-answer pairs without the intermediate reasoning steps used during training. This absence of training sequences makes detection substantially more challenging. (2) LRMs primarily acquire reasoning abilities during two stages: SFT and RL. This potentially provides developers with opportunities to manipulate leaderboard performance by strategically contaminating benchmarks in the earlier stage, while evading detection methods through subsequent training. Given these challenges, the effectiveness of existing detection methods against LRM contamination remains uncertain.
Figure 1: Two scenarios where contamination may happen to LRMs. In Stage I (pre-LRM), while SFT contamination to the base model is initially detectable, contamination evidence can be concealed through subsequent RL training. In Stage II (post-LRM), extensive contamination with CoT on advanced LRMs barely leaves evidence for existing memorization-based detection methods.
In this paper, we present the first systematic study of benchmark contamination in LRMs, structured around two points where contamination can happen. In particular, Stage I (pre-LRM) investigates contamination introduced to the base model while acquiring reasoning ability via SFT and RL; Stage II (post-LRM) investigates contamination applied to an advanced LRM as a final SFT step. Under each stage, we comprehensively evaluate the effectiveness of existing detection methods.
Stage I (pre-LRM): contamination happens when the base model evolves into LRMs. We simulate contamination introduced during the period which the base model acquires reasoning ability through SFT and RL. After evaluating 10 representative contamination detection methods spanning generation-based, perturbation-based, reference-based, and reference-free approaches, we find that while SFT contamination to the base model is initially detectable, contamination evidence can be concealed through subsequent GRPO (Shao et al., 2024) training with clean samples. To isolate the core reasons behind GRPO’s ability to conceal contamination, we conducted carefully designed controlled experiments to rule out the possibility that simply training with more clean samples results in the observed concealment, pointing to the conclusion that the GRPO optimization objective might be the primary driver for obscuring contamination. Then, we performed a theoretical analysis showing that the PPO-style importance sampling/clipping gate can drive the drop in detection performance. Our ablation studies confirm that while plain rejection sampling (RAFT) will not shrink the member/non-member separability, its variant RAFT++ (Xiong et al., 2025) that adds on the importance sampling/clipping term again makes detection harder. As many RL algorithms adopt similar training objectives, this demonstrates a significant risk to the integrity of benchmark evaluations.
Stage II (post-LRM): contamination with CoT applied to LRMs. We simulate contamination with CoT introduced to advanced LRMs as the final training step. Surprisingly, although exclusively SFT on the benchmark samples with CoT yields a huge inflated performance, it leaves little evidence to existing detection approaches: almost all the detection approaches consistently perform near random guess in all the benchmarks. The log-prob distributions of both members and non-members show that without exposure to non-members, contaminated LRMs still have more confidence when responding to those unseen samples that are similar to the training set. This may undermine the key assumption behind many existing detection techniques that the benchmark contamination problem is primarily about memorizing samples (Morris et al., 2025; Hayes et al., 2025).
Overall, our findings reveal that existing contamination detection methods are fragile under LRM contamination scenarios: RL conceals SFT contamination evidence introduced during the transition from base models to LRMs, while contamination with CoT applied to advanced LRMs leaves little detectable evidence. These findings underscore the urgent need for advanced contamination detection methods and trustworthy evaluation protocols tailored to LRMs. Accordingly, we outline potential directions for guaranteeing the integrity of evaluating LRMs (Section 5). We hope that our discoveries will inspire further research dedicated to building fair evaluation arenas for LRMs.
2 Related Works
LRMs. LRMs achieve superior performance on challenging mathematical and coding tasks (Team et al., 2025), driven by inference-time scaling (Jaech et al., 2024; Snell et al., 2024; Zhang et al., 2025). To endow reasoning abilities to existing models, numerous efforts have been focusing on either SFT distillation (Li et al., 2025; Muennighoff et al., 2025; Guha et al., 2025; Ye et al., 2025; Bercovich et al., 2025) or RL with verifiable rewards (Liu et al., 2025a; Zeng et al., 2025; Yue et al., 2025). In SFT distillation, model developers distill knowledge from advanced LRMs into smaller models (Guo et al., 2025). While RL enables models to generate rollouts and receive rewards from verifiers, improving models’ reasoning ability through feedback (Liu et al., 2025a; Zeng et al., 2025; Yue et al., 2025; Liu et al., 2025b). These two stages create many opportunities for developers to contaminate the benchmarks and evade detection.
Benchmark Contamination Detections. Benchmark contamination detection methods aim to identify whether evaluation datasets have been exposed during training (Oren et al., 2023). Prior work has proposed approaches based on: instance similarity (Karamolegkou et al., 2023), probability analysis (Mattern et al., 2023), instance generation (Deng et al., 2023; Ranaldi et al., 2024), and answer memorization (Yim et al., 2024). In this work, we select representative methods applicable to our setting, from probability analysis and instance generation, and further categorize them into: generation-based (Dong et al., 2024; Wu et al., 2025), perturbation-based (Li et al., 2025; Mattern et al., 2023), reference-based (Mireshghallah et al., 2022; Carlini et al., 2021), embedding-based (Tu et al., 2024; Liu et al., 2024), and reference-free (Zhang et al., 2024; Li et al., 2025; Yeom et al., 2018; Shi et al., 2023) methods. Each of these relies on distinct assumptions (Fu et al., 2024), and their effectiveness in the LRMs contamination scenario remains underexplored.
Benchmark Contamination Concealment. Model developers hope to conceal contamination evidence while still having performance inflation. Prior work has explored evading detection through benchmark augmentation, such as rephrasing solutions with strong LLMs (Dekoninck et al., 2024; Samuel et al., 2024), but in LRM settings, most benchmarks only have question–answer pairs without step-by-step solutions, making such methods inapplicable. (Bordt et al., 2024) explores from the training dynamic perspective, showing that performance inflation due to contamination diminishes as pre-training progresses. To our knowledge, we are the first to investigate contamination concealment at the algorithmic level.
3 RL Conceals Contamination (Stage I: pre-LRM)
Contamination Setup. We define SFT contamination as the model being exposed to both the benchmark question and responses distilled from an advanced LRM, where RL contamination refers to the model encountering the benchmark question and having received rewards based on its generated responses during RL finetuning. For each dataset, we randomly sample half of the questions as the member set (used for contamination) and leave the remaining half as the non-member set (for detection evaluation). More details about our contamination pipelines, datasets, and implementation can be found in appendix D.1,  D.3, and  D.4.
Detection Setup. We consider 10 representative detection methods. For each question, we generate 8 responses and compute the detection value on each response, then average these values to obtain a final detection score for the question. For the rationale and ablation studies of choosing responses to compute the detection scores, please refer to Appendix E.1. We report Area Under the Receiver Operating Characteristic (AUROC) by comparing detection scores between member and non-member sets within the same benchmark. Higher AUROC values indicate better detection.
3.1 GRPO Conceals Benchmark Contamination
Contamination Inflation Mainly Comes From SFT. We evaluate multiple contamination scenarios that may happen during SFT and RL and summarize the empirical results in Tab. 1. Results show that clean SFT training yields an 11.30% improvement in pass@1 performance, while SFT contamination further inflates results by an additional 8.82% on average across six benchmarks. In contrast, RL contamination, despite introducing the benchmark questions and giving rewards based on the model-generated responses, shows no significant difference compared to using a clean RL training set. These findings indicate that after extensive SFT training, introducing benchmark samples during the RL process provides little performance benefit.
Table 1: Pass@1 (%) under different contamination scenarios when the base model evolves into LRMs. Empirical results demonstrate that contamination inflation mainly comes from the SFT stage. “/” means not used, and “Mem” denotes members. We first train the base model with SFT and then RL. The last row denotes the results of the base model.
SFT Data RL Data Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg.
Clean & Mem Clean & Mem 52.56 44.70 30.00 30.00 39.52 73.00 44.96
Clean & Mem Clean 52.52 45.71 34.67 28.00 39.89 72.50 45.55
Clean & Mem / 53.77 49.58 31.62 32.73 40.74 74.92 47.23
Clean Clean & Mem 44.62 40.74 24.85 27.88 35.23 65.00 39.72
Clean Clean 47.11 41.41 24.44 26.67 32.72 70.83 40.53
Clean / 44.35 40.34 24.79 23.54 34.24 63.20 38.41
/ / 36.48 32.20 2.50 10.83 28.58 52.50 27.18
Table 2: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Higher AUROC, better detection performance. Each AUROC is averaged over detection scores from 8 rollouts.
Contamination Detection Methods Training Stages Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Generation based
Verbatim (Wu et al., 2025) Before RL 47.58 49.86 47.56 53.56 52.52 65.50 52.76 +0.00
RL w/ Clean 45.60 51.28 47.56 56.44 52.05 60.00 52.16 -0.60
RL w/ Clean&Mem 46.17 50.34 52.67 55.56 51.71 63.62 53.35 +0.59
CDD (Dong et al., 2024) Before RL 55.75 57.32 41.56 59.11 59.27 61.75 55.80 +0.00
RL w/ Clean 55.47 51.08 43.33 60.00 60.18 62.00 55.34 -0.46
RL w/ Clean&Mem 56.32 44.14 35.56 65.11 60.31 49.38 51.80 -3.95
Perturbation based
Neighbor (Mattern et al., 2023) Before RL 54.76 41.19 50.00 41.56 55.64 61.10 50.71 +0.00
RL w/ Clean 54.10 39.68 50.67 44.22 53.42 60.50 50.43 -0.28
RL w/ Clean&Mem 53.05 41.08 50.44 52.67 68.16 64.00 54.90 +4.19
Reference based
LiRA (Mireshghallah et al., 2022) Before RL 85.37 86.80 100.00 82.00 87.01 93.62 89.13 +0.00
RL w/ Clean 74.41 84.65 70.22 87.78 81.04 82.75 80.14 -8.99
RL w/ Clean&Mem 69.73 77.85 63.11 82.22 79.05 77.38 74.89 -14.24
Ref (Carlini et al., 2021) Before RL 73.27 63.30 60.22 41.11 73.10 82.00 65.50 +0.00
RL w/ Clean 66.77 58.41 45.33 51.11 65.54 73.62 58.08 -7.42
RL w/ Clean&Mem 62.77 54.17 43.11 50.44 65.38 72.62 58.86 -6.64
Reference free
Zlib (Carlini et al., 2021) Before RL 49.38 58.61 73.56 43.56 50.19 45.00 53.38 +0.00
RL w/ Clean 45.94 54.99 66.22 35.56 46.65 39.38 48.12 -5.26
RL w/ Clean&Mem 46.04 55.30 64.89 28.89 44.87 39.00 44.74 -8.64
Min–K%++ (Zhang et al., 2024) Before RL 47.57 50.90 41.90 59.11 52.27 45.88 49.61 +0.00
RL w/ Clean 46.25 46.78 36.67 50.89 51.35 29.62 43.59 -6.02
RL w/ Clean&Mem 43.77 48.21 21.78 38.00 48.91 43.62 40.72 -8.89
Min–K% (Shi et al., 2023) Before RL 69.19 69.51 85.56 75.56 71.16 78.75 74.96 +0.00
RL w/ Clean 55.19 60.60 62.89 65.56 61.50 61.87 61.27 -13.69
RL w/ Clean&Mem 53.93 59.74 59.56 62.67 57.31 59.25 58.54 -16.42
Max–K% (Maini et al., 2024) Before RL 64.50 64.31 65.11 81.78 67.27 76.00 69.83 +0.00
RL w/ Clean 53.05 51.43 49.78 50.22 51.84 57.75 52.35 -17.48
RL w/ Clean&Mem 49.03 51.04 50.00 50.00 52.34 47.50 49.99 -19.84
Loss (Carlini et al., 2021) Before RL 69.18 69.81 86.22 77.33 70.95 79.38 75.48 +0.00
RL w/ Clean 55.22 60.50 62.44 65.78 61.50 62.12 61.26 -14.22
RL w/ Clean&Mem 53.99 60.01 59.33 62.67 57.40 59.38 58.80 -16.68
To understand whether current contamination detection methods can still successfully detect contamination in LRMs, and whether RL training can alter the signals exploited by contamination detectors, we evaluate SFT-contaminated models before and after GRPO. Tab. 2 reveals systematic shifts in AUROC across diverse detection methods. Our analysis highlights three key observations:
SFT contamination can be detectable at first. Several reference-free approaches (Min-K% (Shi et al., 2023), Max-K% (Maini et al., 2024), and LOSS (Carlini et al., 2021)) can detect SFT contamination at a certain level, achieving AUROC around 73.42% across six contaminated benchmarks. The reference-based detection approach, LiRA (Mireshghallah et al., 2022), which assumes access to the training data distribution, also demonstrates superior performance with an average AUROC of 89.13% across six benchmarks.
GRPO conceals contamination. After applying GRPO to the SFT-contaminated model, we observe a consistent decrease in AUROC across all detection methods and benchmarks. We further analyze the average log probability of member and non-member samples before and after GRPO training. Fig.3 shows two key patterns: (1) GRPO lowers the entropy of generated sequences, indicating that the model becomes more confident in its generation, which is consistent with prior observations in (Cui et al., 2025); (2) the log probability distribution of members and non-members converge after GRPO. Since the gaps in log probability are the core statistical backbone of existing contamination detectors, these findings suggest that GRPO may inherently suppress contamination evidence by rendering members and non-members indistinguishable.
[…]
(a) Results on GPQA-Diamond
(b) Results on OlympiadBench
Figure 3: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training. After further GRPO with clean samples on the SFT contaminated model, the log-prob distributions of members and non-members become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. More log-prob distributions can be found in Fig. 56, and 7.
More GRPO, less contamination evidence. To examine whether the concealment effect strengthens with additional training, we extend GRPO to SFT-contaminated models using 10K questions from DeepMath-103K (He et al., 2025) for one epoch (156 steps). As shown in Fig.2, AUROC consistently decreases across all detection methods and benchmarks as the number of GRPO steps increases. Given that our maximum 156 training steps are still far fewer than the steps used in some advanced open-sourced reasoning models (Luo et al., 2025b; a), we expect that extensive GRPO training would render all existing detection methods to near-random performance eventually.
Further training will not make models forget contamination. One possible explanation is that additional training makes models forget contamination, thus detections perform random guessing and pass@1 match the clean SFT baseline. To test this, we examine it with two experiments. First, we train SFT contaminated models with GRPO on both clean and contaminated datasets. As shown in Tab. 2, we observe a comparable drop in AUROC relative to the no RL baseline, similar to performing RL solely on clean data. Also, the contaminated model, further trained with GRPO, still shows an average performance inflation of 7.14% across six benchmarks and does not fall back as the clean SFT model, shown in Tab. 1. Second, we continue SFT on the SFT contaminated model with an additional 4 epochs on clean data. Fig. 2 and Tab. 14 demonstrate that further SFT is unable to conceal the benchmark contamination, while the pass@1 would continue to rise. Together, these results show that subsequent GRPO training preserves performance inflation while reducing detectable evidence of contamination may have some underlying reasons, rather than simply forgetting the contamination after further training.
3.2 Theoretic Analysis
In this section, we perform theoretical analysis to demonstrate that PPO-style clipping and importance sampling are the root cause of the concealment. Intuitively, the importance sampling and clipping term reweights terms so that the most off-policy trajectories are damped by the clip while typical on-policy ones keep their influence. This reweighting hits non-members more as they have more extreme successes, so clipping cuts misaligned influence and lets ordinary, on-policy successes steer the update. With more headroom, non-member’s NLL drops more and the gap contracts.
Setup. We denote $\ell(x,y)$ to be the negative log likelihood (NLL) of the current model of generating $y$ given prompt $x$, members as $M$ and non-members as $N$, policy model $\pi_{k}$ at step $k$. We focus on analyzing the gap $G_{k}$ of negative log likelihood for members and non-members on correct samples (i.e., $r=1$), as assessing contamination on erroneous outputs is not especially meaningful. Formally, we can write
$G_{k}=\mathbb{E}_{x\sim N}\mathbb{E}_{y\sim\pi_{k}(\cdot|x)}[\ell_{k}(x,y)\mid r=1]$ $-\mathbb{E}_{x\sim M}\mathbb{E}_{y\sim\pi_{k}(\cdot|x)}[\ell_{k}(x,y)\mid r=1]$ (1)
If this gap contracts, i.e., $G_{k+1}-G_{k}<0$, members and non-members become closer in the NLL sense, making contamination detection harder since many methods (Zhang et al., 2024; Shi et al., 2023; Maini et al., 2024; Carlini et al., 2021) are based on the separation of NLLs. For a fixed prompt $x$, we define the NLL drift as
$\Delta_{x}:=\mathbb{E}_{\pi_{k+1}}[\ell_{k+1}\mid r=1,x]-\mathbb{E}_{\pi_{k}}[\ell_{k}\mid r=1,x].$ (2)
We notice that we can rewrite the NLL gap as
$G_{k+1}-G_{k}:=\mathbb{E}_{x\in N}[\Delta_{x}]-\mathbb{E}_{x\in M}[\Delta_{x}].$ (3)
In our following analysis, we thus focus on investigating the behavior of $\Delta_{x}$ on members and non-members. If an algorithm yields on average smaller $\Delta_{x}$ on non-members, the algorithm should be able to conceal contamination.
Notations. At token $t$, let $A_{t}$ be the method’s per token reward/advantage and $w_{t}$ the weight from importance sampling and clipping. Define
$A_{t}^{w}:=w_{t}A_{t},\quad\bar{A}^{w}(s):=\mathbb{E}_{a\sim\pi_{k}(\cdot\mid s)}[A^{w}(s,a)],\quad\tilde{A}_{t}^{w}:=A_{t}^{w}-\bar{A}^{w}(s_{t})$ (4)
to measure how good a state is compared to the average. In particular, $w_{t}\!=\!\rho_{t}m_{t}$ with $\rho_{t}\!=\!\pi_{\theta}(a_{t}\mid s_{t})/\pi_{\text{old}}(a_{t}\mid s_{t})$ being the importance sampling and $m_{t}\in\{0,1\}$ be a mask indicating if the clipping is activated, specifically $m_{t}\!=\!0$ indicates that there is no gradients from the update. Moreover, we define $p_{k}(x)\!=\!\mathbb{E}_{y\sim\pi_{k}(\cdot\mid x)}[r(x,y)]$ to be the overall success rate of the prompt, and a value function as $q_{k}(s,a)\!:=\!\mathrm{Pr}(r=1|s,a)$ and $p_{k}(s)\!:=\!\mathbb{E}_{a\sim\pi_{k}(\cdot\mid s)}[q_{k}(s,a)]$ for success rate at that state. And we define $B(s)\!=\!\mathbb{E}_{a\sim\pi}[\rho(s,a)m(s,a)q_{k}(s,a)]$ and $C(s)\!=\!\mathbb{E}_{a\sim\pi}[\rho(s,a)m(s,a)]$. We assume that the RL training is performed on the benchmark data (i.e., training data is the combination of members $M$ and non-members $N$), and it is in a tabular setting for simplicity. Since members have been utilized during training, it is natural to assume the $p_{k}(s)$ for members are larger than non-members, and the NLL for members is lower than non-members.
Table 3: AUROC (%) of detection approach, Loss (Carlini et al., 2021), evaluated on SFT contaminated model further trained with different RL objectives. The gray row indicates no ablation on the objective, and ✗ means remove the term from the objective. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). RL steps are 64, or the step before the model collapses. The results show that clipping is the main driver for the contraction, which aligns with our theory.
Training
Objectives
Clipping
Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
[…]
RAFT++ 69.16 73.68 74.44 76.22 71.08 81.75 74.39 -1.09
[…]
GRPO 68.83 70.20 80.44 73.78 68.30 78.12 73.28 -2.20
Theorem 3.1.
For a small natural gradient step with step size $\eta$ on a PPO style loss, we have
$\Delta_{x}$ $=-\eta\,\underbrace{\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\tilde{A}_{t}^{w}\,\middle|\,r=1,x\right]}_{\text{(A) $\mu(x)$}}+\eta\,\underbrace{\operatorname{Cov}\!\left(\ell_{k},\;\sum_{t=1}^{T}\tilde{A}_{t}^{w}\right)}_{\text{(B) covariance $\beta(x)$}}+O(\eta^{2})\,$ (5)
The proof can be found in appendix C. Intuitively, $\mu(x)$ measures the average push on the example’s NLL from correct trajectories, where $\beta$ serves as a reweighting term accouting for the importance sampling/clipping. Here we consider several instantiations using different algorithms to investigate the core driver for contraction. The training objectives for each algorithm are listed in Appendix B.
RAFT. In plain rejection sampling, we have $w_{t}\!=\!1$ and $A_{t}\!=\!\mathbf{1}\{r=1\}$, so on correct trajectories
$\tilde{A}_{t}^{w}=1-p_{k}(s_{t}),\quad\mu^{\mathrm{RAFT}}(x)=\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\big(1-p_{k}(s_{t})\big)\,\middle|\,r=1,x\right].$
The covariance term is
$\beta^{\mathrm{RAFT}}(x)=\operatorname{Cov}\!\Big(\ell_{k},\;\sum(1-p_{k}(s_{t}))\Big)=-\operatorname{Cov}\!\Big(\ell_{k},\;\sum p_{k}(s_{t})\Big).$
We note that lower loss $\ell_{k}$ corresponds to higher probabilities $p_{k}(s_{t})$, and thus $\beta^{\mathrm{RAFT}}(x)\!>\!0$. Moreover, non-members correct trajectories can exhibit much higher variance in loss and probabilities, thus, the $\beta_{N}$ term is typically larger than $\beta_{M}$. Consequently,
$\Delta_{N}-\Delta_{M}=-\eta\big(\mu_{N}-\mu_{M}\big)+\eta\big(\beta_{N}-\beta_{M}\big),$
where both gaps $(\mu_{N}-\mu_{M})$ and $(\beta_{N}-\beta_{M})$ are positive. Empirically, the covariance gap offsets the mean gap, yielding $\Delta_{N}-\Delta_{M}\geq 0$, i.e., RAFT is unable to conceal contamination evidence.
RAFT++. Using the same $A_{t}=\mathbf{1}\{r=1\}$, on $r=1$ paths
$\tilde{A}_{t}^{w}=\rho_{t}m_{t}-B_{k}(s_{t}),\quad\mu^{\mathrm{RAFT{++}}}(x)=\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\big(\rho_{t}m_{t}-B_{k}(s_{t})\big)\,\middle|\,r=1,x\right].$
We note that the difference of $\mu$ cannot possibly lead to large deviations between members/non-members as $0\!\leq\rho_{t}m_{t}\!\leq\!1+\epsilon\!$ and $B_{k}(s_{t})\!\leq\!1$ for both groups and the term is normalized by length. For the covariance term though, we have
$\beta^{\mathrm{RAFT{++}}}(x)=\operatorname{Cov}\!\Big(\ell_{k},\;\sum(\rho_{t}m_{t}-B_{k}(s_{t}))\Big)=\operatorname{Cov}\!\Big(\ell_{k},\;\sum\rho_{t}m_{t}\Big)\;-\;\operatorname{Cov}\!\Big(\ell_{k},\;\sum B_{k}(s_{t})\Big).$
Compared to RAFT, the new term $\operatorname{Cov}(\ell_{k},\sum\rho_{t}m_{t})$ is negative as correct path with higher loss are anomaly and typically got clipped more. Moreover, this is much more prominent in non-members due to high variance in correct trajectories loss. The second covariance term, although still negative, are not that significant for non-members compared to members due to an average over all possible actions. Therefore, overall it leads to
$\Delta_{N}-\Delta_{M}=-\eta(\mu_{N}-\mu_{M})+\eta(\beta_{N}-\beta_{M})\;<\;0,$
i.e., RAFT++ contracts the membership gap. The driver is precisely the PPO-style importance sampling/clipping: it removes the RAFT covariance cancellation by making $\operatorname{Cov}(\ell_{k},\sum\rho m)$ non-positive and more negative for non-members.
GRPO. Finally, we investigate the GRPO contraction term. To ease the analysis, we consider an idealized setting where we define the advantage term as $A_{k}(x,y)\!=\!r(x,y)\!-\!p_{k}(x)$ with no standard deviation term and $\tilde{A}_{t}^{w}\!=\!\tilde{A}_{t}^{\text{RAFT}}\!-\!p_{k}(x)(\rho_{t}m_{t}\!-\!C(s_{t}))$. Clearly, we have
$\mu^{\text{GRPO}}(x)$ $=\mu^{\text{RAFT++}}(x)-p_{k}(x)\mathbb{E}\!\left[\frac{1}{T}\sum\big(\rho_{t}m_{t}-C_{k}(s_{t})\big)\,\middle|\,r=1,x\right]$
$\beta^{\text{GRPO}}(x)$ $=\beta^{\text{RAFT++}}(x)-p_{k}(x)\operatorname{Cov}\!\Big(\ell_{k},\;\sum(\rho_{t}m_{t}-C_{k}(s_{t}))\Big)$
By similar argument, we know that the $\mu$ term does not contribute significantly to the concealment. The covariance term can be analyzed similarly to show that the concealment also happen on GRPO thanks to the importance sampling and clipping term.
3.2.1 Empirical Support
To confirm empirically the prediction of our theoretical results, we evaluate the Loss detector (Carlini et al., 2021) after training with RAFT (Dong et al., 2023)/RAFT++ (Xiong et al., 2025)/GRPO. The overall results can be found in table 3. From the results, there are several observations.
Effect on detectability. Under RAFT, the Loss detector (Carlini et al., 2021) performance remains essentially unchanged relative to the SFT contaminated baseline w/o further RL. In contrast, RAFT++ and GRPO (with clipping enabled) produce a sharp drop in detector performance.
Importance sampling vs. clipping. The clipping term, often treated purely as a training stabilizer, materially contributes to concealment, as predicted by theory. When we retain importance sampling but remove clipping in RAFT++ and GRPO, both algorithms show little to no reduction in Loss-detector performance (Table 3). Intuitively, as the clip threshold $\epsilon\!\to\!\infty$, the effective weight satisfies $\sum_{t}\rho_{t}m_{t}\approx T$, and the covariance term in our decomposition tends toward zero for both members and non-members, eliminating the shrinkage effect.
These two observations perfectly reflect our theoretical analysis, empirically validating that the PPO-style importance sampling/clipping term is the key driver behind GRPO contamination concealment. Given that many RL algorithms adopt this term in their objectives, this suggests that a broad class of RL methods may inherently exhibit similar concealment capability.
4 Contamination with CoT on advanced LRMs Barely Leaves Evidence (Stage II: post-LRM)
Contamination Setup. In this setup, we simulate contamination with CoT applied to advanced LRMs at the final stage of training. We use DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), and checkpoints from OpenThought3 (Guha et al., 2025) as the initial models. We simulate extensive contamination with CoT by applying SFT exclusively on the member data in this section. Additional implementation details are provided in Appendix D.4.
Tab. 4 and 5 show the results of pass@1 on six reasoning benchmarks and AUROC of detection approaches performance (w/ the same detection setup as Stage I), respectively. We observe that:
Extensive SFT Contamination with CoT results in a huge performance inflation. As shown in Tab. 4, LRMs can substantially benefit from extensive contamination with CoT. For instance, DeepSeek-R1-Distill-Llama-8B model exhibits an average performance inflation of up to 11.76% across six benchmarks after being exposed to membership data seven times. Such inflation enables contaminated LRMs to artificially boost performance in benchmarks and have an overrated rank in the reasoning leaderboard with little extra training cost.
Table 4: Pass@1 (%) of advanced LRMs before and after SFT contamination with CoT.
Models Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg.
DeepSeek-R1-Distill-Llama-8B 52.10 43.94 33.33 43.33 32.97 84.58 48.38
$\hookrightarrow$ w/ extensive SFT Contamination 61.83 53.16 51.67 61.67 38.74 93.75 60.14
DeepSeek-R1-Distill-Qwen-7B 55.70 48.65 39.26 53.70 37.25 91.94 54.42
$\hookrightarrow$ w/ extensive SFT Contamination 58.77 50.87 42.59 58.91 40.81 90.67 57.10
OpenThinker3-7B (15K) 50.81 41.67 21.67 29.17 34.01 77.50 42.47
$\hookrightarrow$ w/ extensive SFT Contamination 52.74 47.64 33.33 30.48 40.56 78.70 47.25
[…]
Contamination Detection Methods Init Models Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg.
Generation based
Verbatim (Wu et al., 2025) DS Llama 48.73 50.45 41.33 61.56 59.10 40.63 50.30
DS Qwen 46.87 55.85 60.44 68.89 56.87 50.63 56.59
OpenThink 43.78 55.36 60.89 56.67 51.78 42.38 51.81
CDD (Dong et al., 2024) DS Llama 51.84 53.83 60.00 53.11 58.08 57.50 55.73
DS Qwen 51.46 48.29 50.00 53.78 54.71 41.00 49.87
OpenThink 49.98 50.23 53.31 51.24 54.52 50.44 51.62
Perturbation based
Neighbor (Mattern et al., 2023) DS Llama 49.94 39.32 53.11 43.33 49.68 60.00 49.23
DS Qwen 52.99 40.29 62.44 49.33 55.34 54.87 52.54
OpenThink 53.76 42.95 34.00 42.22 52.89 51.50 46.22
Reference based
LiRA (Mireshghallah et al., 2022) DS Llama 57.92 53.01 53.56 75.33 69.44 58.75 61.34
DS Qwen 46.52 43.93 50.22 58.89 59.33 54.00 52.15
OpenThink 62.35 64.77 58.44 64.44 64.81 61.62 62.74
Ref (Carlini et al., 2021) DS Llama 53.79 46.50 46.44 64.00 63.57 51.25 54.26
DS Qwen 53.30 44.37 46.89 44.22 53.09 41.75 47.27
OpenThink 57.34 49.86 37.56 50.44 59.30 69.12 53.94
Reference free
Zlib (Carlini et al., 2021) DS Llama 49.52 54.74 64.22 37.11 45.97 47.12 49.78
DS Qwen 46.52 57.38 64.89 36.89 43.30 42.12 48.52
OpenThink 45.65 55.37 74.22 36.89 43.51 36.62 48.71
Min–K%++ (Zhang et al., 2024) DS Llama 55.45 59.10 45.95 70.22 60.89 57.50 58.19
DS Qwen 48.92 56.83 48.44 59.33 51.83 62.62 54.66
OpenThink 51.85 58.31 66.44 55.00 49.41 41.05 53.68
Min–K% (Shi et al., 2023) DS Llama 57.86 61.68 53.33 72.67 67.12 61.87 62.42
DS Qwen 49.75 53.93 51.78 61.56 54.50 56.75 54.71
OpenThink 53.52 57.19 60.44 57.56 54.83 47.37 55.15
Max–K% (Maini et al., 2024) DS Llama 53.85 55.96 50.67 60.44 59.22 52.50 55.44
DS Qwen 49.65 50.92 40.44 73.33 54.08 56.25 54.11
OpenThink 55.12 58.29 46.22 79.33 54.20 59.38 58.76
Loss (Carlini et al., 2021) DS Llama 57.91 61.78 52.89 73.56 67.00 62.38 62.59
DS Qwen 49.77 54.09 52.00 63.78 54.76 56.75 55.19
OpenThink 53.44 57.61 61.33 56.67 55.07 48.12 55.37
Extensive contamination with CoT on LRMs barely leaves evidence. As illustrated in Tab. 5, detection methods, which were effective in contamination introduced when the base model evolves into LRMs, consistently fail under extensive contamination with CoT to LRMs, performing close to random guessing. The previous SOTA approach, LiRA (Mireshghallah et al., 2022), achieves only 58.74% AUROC on average across six benchmarks. Then, we analyze the log prob of member and non-member samples before and after final stage contamination, shown as Fig. 4. After the extensive SFT contamination with CoT on members, the log prob of both members and non-members increases at a similar margin. This indicates that even without exposure to non-members, contaminated LRMs still have more confidence when responding to unseen samples that share similar distributions to training samples, which also explains why extensive contamination with CoT on LRMs barely leaves evidence. These results suggest that model developers could extensively contaminate their LRMs in the final stage while leaving little detectable evidence.
(a) R1 Distill LLaMA results on OlympiadBench
(b) R1 Distill Qwen results on OlympiadBench
Figure 4: Log-prob distributions for members vs. non-members of advanced LRMs before and after SFT contamination. After extensive SFT contamination on members, the log prob of both members and non-members increases at a similar margin. More figures are in Fig. 8 and 9.
Discussion. Despite Feng et al. (2024) demonstrating that contamination detection could work in non-reasoning model scenarios, the detectors do not have access to the reasoning trajectories used in the LRM contamination scenario, so they have to rely on the generated responses from LRMs. However, LRMs typically possess strong reasoning abilities to output step-by-step long CoT and are difficult to converge on a specific sequence after contamination with long CoT. This may indicate that rather than memorizing specific reasoning trajectories, LRMs internalize the underlying knowledge and reasoning process during the contamination with CoT data, enabling generalization to distributionally similar questions (e.g., non-members). However, existing detection approaches typically assume that contaminated models assign lower loss to training sequences than to unseen ones. So, due to this generalization ability, LRMs can also achieve low loss on unseen samples that resemble the training set. This fundamentally challenges the assumption that all the detection approaches rely on, which is that benchmark contamination is more about memorizing the benchmark samples (Morris et al., 2025; Hayes et al., 2025).
5 Conclusion
We present the first systematic study of benchmark contamination in LRMs, structured around two points where contamination can happen. Our results reveal a critical vulnerability in LRM evaluation: contamination detection methods are fragile and contamination introduced at either stage can be concealed. In Stage I (pre-LRM), while SFT contamination to the base model is initially detectable, contamination evidence can be concealed through subsequent RL training. In Stage II (post-LRM), extensive contamination with CoT on advanced LRMs barely leaves evidence for existing memorization-driven detection methods. Our findings call for an urgent need of protocols that ensure fair evaluations among LRMs. Here, we propose two potential directions to ensure it: (I) Model developers should release more intermediate training checkpoints, enabling the community to better monitor and regulate potential benchmark contamination in each training stage. (II) Researchers working on contamination detections should advance beyond memorization-driven methods and explicitly account for the long CoT reasoning and generalization capacity of LRMs. Despite the assumption that contamination is about memorizing the training data inspires numerous detection methods before the LRM era, it may become outdated right now. Detection approaches that are solely based on log-probs or mitigation approaches such as minor benchmark modifications, are definitely inadequate in this context and risk systematically failing. These findings all highlight the need for new assumptions in contamination detection and the development of contamination-robust evaluation protocols in the LRM setting.
Appendix A Limitations
Our work reveals critical vulnerabilities: RL fine-tuning can conceal benchmark contamination when base models evolve into LRMs; contamination with CoT applied to advanced LRMs leaves little evidence detectable by existing memorization-based methods. Although we do not propose a new detection algorithm to mitigate these risks, we reveal that the failure of current detection approaches stems from their reliance on log-probability and on the assumption that training samples consistently incur lower loss than unseen samples. Given the unique characteristics of LRMs, future detection methods must adopt more advanced assumptions to address this fundamental challenge. By highlighting these risks, we aim to spur further research on robust and trustworthy evaluation protocols for LRMs.
Appendix B Algorithms
SFT
Let $\mathcal{D}=\{(q,o)\}$ be a corpus of questions $q$ and responses $o$, where $o$ are distilled from advanced LRMs. Let $\pi_{\theta}(o\mid q)$ be an autoregressive policy model. The $\pi_{\theta}$ is then fine-tuned to maximize the log-likelihood over the responses:
$\mathcal{L}_{\text{SFT}}(\theta)$ $=\mathbb{E}_{(q,o)\sim\mathcal{D}}\left[-\log\pi_{\theta}(o\mid q)\right]$
GRPO
For each question $q$, GRPO samples a group of outputs $\{o_{1},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{L}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{\,q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(O\mid q)}$
$\Big[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Big(\min\!\Big(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)}\,A_{i},\;\operatorname{clip}\!\Big(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)},\,1-\varepsilon,\,1+\varepsilon\Big)\,A_{i}\Big)-\beta\,\mathrm{D}_{\mathrm{KL}}\!\left(\pi_{\theta}\middle\|\pi_{\mathrm{ref}}\right)\Big)\Big],$
$\mathrm{D}_{\mathrm{KL}}\!\left(\pi_{\theta}\middle\|\pi_{\mathrm{ref}}\right)=\frac{\pi_{\mathrm{ref}}(o_{i}\mid q)}{\pi_{\theta}(o_{i}\mid q)}-\log\!\frac{\pi_{\mathrm{ref}}(o_{i}\mid q)}{\pi_{\theta}(o_{i}\mid q)}-1,$
where $\varepsilon$ and $\beta$ are hyper-parameters. We denote $r\in\{0,1\}$ as a binary reward function that assigns scalar feedback to a question-output pair. $A_{i}$ is the advantage, computed using a group of rewards $\{r_{1},r_{2},\ldots,r_{G}\}$ corresponding to the outputs within each group:
$A_{i}\;=\;\frac{r_{i}-\operatorname{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\operatorname{std}(\{r_{1},r_{2},\ldots,r_{G}\})}.$
RAFT
The RAFT is also referred to as the rejection-sampling fine-tuning. The algorithm consists of two parts:
  • Dataset Collection. For each question $q$, we sample a group of outputs $\{o_{1},\cdots,o_{G}\}$. For each response $o_{i}$, we compute the reward $r_{i}\in\{0,1\}$. We then retain only the responses that receive a reward of 11 and put them in a dataset $\mathcal{D}$.
  • Model Fine-tuning. The current policy $\pi_{\theta}$ is then fine-tuned to maximize the log-likelihood over the selected dataset:
    $\mathcal{L}_{\text{RAFT}}(\theta)$ $=\mathbb{E}_{(q,o)\sim\mathcal{D}}[-\log\pi_{\theta}(o\mid q)]$
RAFT++
RAFT++ is a variant of plain RAFT that also introduced importance sampling and clipping. The loss is very similar to GRPO with no KL divergence term except for the advantage calculation. The advantage for RAFT++ will be:
$A_{i}=\mathcal{I}(r_{i}=1),$
which indicates essentially that we train only on the positive samples.
Appendix C Proof
Here we restate our theorem 3.1, and provide a full proof.
Theorem.
For a small natural gradient step with step size $\eta$ on a PPO style loss, we have
$\Delta_{x}$ $=-\eta\,\mathbb{E}\!\left[\frac{1}{T}\sum_{t=1}^{T}\tilde{A}_{t}^{w}\,\middle|\,r=1,x\right]+\eta\,\operatorname{Cov}\!\left(\ell_{k},\;\sum_{t=1}^{T}\tilde{A}_{t}^{w}\right)+O(\eta^{2})\,$ (6)
Proof.
Denote $\operatorname{clip}_{\epsilon}(\rho)=\min(\max(\rho,1-\epsilon),1+\epsilon)$. Consider the optimization problem
$\mathcal{L}(\theta)=\mathbb{E}\left[\sum_{t}\min\left(\rho_{t}(\theta)A_{t},\mathrm{clip}_{\epsilon}(\rho_{t}(\theta))A_{t}\right)\right]$ (7)
The gradient of this objective is
$\nabla_{\theta}\min\left(\rho_{t}(\theta)A_{t},\mathrm{clip}_{\epsilon}(\rho_{t}(\theta))A_{t}\right)$ $=m_{t}A_{t}\rho_{t}\nabla_{\theta}\log\pi_{\theta}(y_{t}\mid s_{t})$ (8)
Let $z_{s,a}$ be the logit parameter for action $a$ at state $s$, so that $\pi_{\theta}(a\mid s)=\mathrm{softmax}(z_{s,\cdot})_{a}.$ Then
$\frac{\partial}{\partial z_{s,a}}\log\pi_{\theta}(y_{t}\mid s_{t})=\mathbf{1}\{s_{t}=s\}\big(\mathbf{1}\{y_{t}=a\}-\pi_{\theta}(a\mid s)\big),$ (9)
and thus
$\frac{\partial\mathcal{L}}{\partial z_{s,a}}=\mathbb{E}\!\left[\sum_{t}\big(\rho_{t}\,m_{t}\,A_{t}\big)\big(\mathbf{1}\{y_{t}=a\}-\pi_{\theta}(a\mid s_{t})\big)\right].$ (10)
It is then clear that after one natural-gradient step,
$z_{s,a}^{k+1}-z_{s,a}^{k}\;=\;\eta\big(A^{w}(s,a)-\bar{A}^{w}(s)\big)=\eta\tilde{A}^{w}(s,a),$ (11)
Using a first-order Taylor expansion of $\log\pi$ and equation 11,
$\log\pi_{k+1}(a\mid s)-\log\pi_{k}(a\mid s)=\eta\tilde{A}^{w}(s,a)+O(\eta^{2}).$ (12)
For a trajectory $y=(y_{1},\dots,y_{T})$,
$\log\frac{\pi_{k+1}(y\mid x)}{\pi_{k}(y\mid x)}$ $=\sum_{t=1}^{T}\big(\log\pi_{k+1}(y_{t}\mid s_{t})-\log\pi_{k}(y_{t}\mid s_{t})\big)=\eta\sum_{t=1}^{T}\tilde{A}^{w}(s_{t},y_{t})+O(\eta^{2}).$ (13)
Define
$\Psi_{x}(y):=\sum_{t=1}^{T}\tilde{A}^{w}(s_{t},y_{t}),\quad S_{k}^{w}(x,y):=\frac{1}{T}\sum_{t=1}^{T}\tilde{A}^{w}(s_{t},y_{t}).$
Then, from $\ell_{k}(x,y)=-\frac{1}{T}\sum_{t}\log\pi_{k}(y_{t}\mid s_{t})$ and equation 12,
$\ell_{k+1}(x,y)-\ell_{k}(x,y)=-\frac{1}{T}\sum_{t=1}^{T}\big(\log\pi_{k+1}-\log\pi_{k}\big)=-\,\eta\,S_{k}^{w}(x,y)+O(\eta^{2}).$ (14)
Moreover,
$R_{x}(y):=\frac{\pi_{k+1}(y\mid x)}{\pi_{k}(y\mid x)}=\exp\!\big(\eta\,\Psi_{x}(y)+O(\eta^{2})\big)=1+\eta\,\Psi_{x}(y)+O(\eta^{2}).$ (15)
Add and subtract $\mathbb{E}_{\pi_{k+1}}[\ell_{k}\mid r{=}1,x]$ in $\Delta_{x}$, we have
$\Delta_{x}$ $=\underbrace{\mathbb{E}_{\pi_{k+1}}\!\big[\ell_{k+1}-\ell_{k}\mid r{=}1,x\big]}_{(A)}\;+\;\underbrace{\Big(\mathbb{E}_{\pi_{k+1}}\![\ell_{k}\mid r{=}1,x]-\mathbb{E}_{\pi_{k}}\![\ell_{k}\mid r{=}1,x]\Big)}_{(B)}.$ (16)
Using equation 14 and a change of measure with $R_{x}$,
$(A)$ $=\frac{\mathbb{E}_{\pi_{k}}\!\big[(\ell_{k+1}-\ell_{k})\,R_{x}\,\big|\,r{=}1,x\big]}{\mathbb{E}_{\pi_{k}}\!\big[R_{x}\,\big|\,r{=}1,x\big]}$
$=\frac{\mathbb{E}_{\pi_{k}}\!\big[(-\eta S_{k}^{w}+O(\eta^{2}))\,(1+\eta\Psi_{x}+O(\eta^{2}))\,\big|\,r{=}1,x\big]}{1+\eta\,\mathbb{E}_{\pi_{k}}[\Psi_{x}\mid r{=}1,x]+O(\eta^{2})}$
$=-\,\eta\mathbb{E}_{\pi_{k}}\!\big[S_{k}^{w}\mid r{=}1,x\big]+O(\eta^{2})$
$=-\,\eta\mathbb{E}_{\pi_{k}}\!\left[\frac{1}{T}\sum_{t=1}^{T}\tilde{A}_{t}^{w}\,\middle|\,r{=}1,x\right].$
where the third equality is simply a first order expansion of division. For term (B), we can compute for any integrable $f$,
$\mathbb{E}_{\pi_{k+1}}[f\mid r{=}1,x]=\frac{\mathbb{E}_{\pi_{k}}[f\,R_{x}\mid r{=}1,x]}{\mathbb{E}_{\pi_{k}}[R_{x}\mid r{=}1,x]}=\mathbb{E}_{\pi_{k}}[f\mid r{=}1,x]+\eta\operatorname{Cov}_{\pi_{k}(\cdot\mid r{=}1,x)}\!\big(f,\Psi_{x}\big)+O(\eta^{2}),$
by expanding numerator and denominator using equation 15. Taking $f=\ell_{k}$ gives
$(B)=\eta\operatorname{Cov}_{\pi_{k}(\cdot\mid r{=}1,x)}\!\big(\ell_{k},\Psi_{x}\big)+O(\eta^{2})=\eta\operatorname{Cov}_{\pi_{k}(\cdot\mid r{=}1,x)}\!\left(\ell_{k},\;\sum_{t=1}^{T}\tilde{A}_{t}^{w}\right).$ (17)
Combining the two part we get our theorem. ∎
Appendix D Experiment Setup
D.1 Contamination Pipelines
SFT Contamination
We randomly select 10K samples from OpenThoughts3 (Guha et al., 2025) to form the clean SFT training set, following the same ratio for each domain as the original paper, to obtain the best results. For the SFT contaminated training set, we use QwQ-32B (Team, 2025) as an advanced LRM to help distillation for the member set. We adopt the temperature 0.6, top-p value of 0.95, and maximum output token with 32768 for the distillation. We use rejection sampling with 64 rollouts, selecting correct trajectories when available. If none of the 64 rollouts produce a correct answer, we randomly select an incorrect trajectory. We replicate the question in the member set with responses 3 times to create the SFT contamination training set. So the training set consists of 11866 samples.
Table 6: Proportion of questions solved after up to 64 rollouts with QwQ-32B.
Olypaidbench GPQA-Diamond AIME2025 AIME2024 Minerva Math AMC2023 Avg.
80.59 92.42 93.33 93.33 56.62 100.00 86.05
RL Contamination
For RL contamination, we replicate the questions in the member set 2 times, and randomly select 4096 samples from DeepMath-103K (He et al., 2025) as the clean RL training set. We choose GRPO as the RL algorithm and train the model with one epoch.
Base Model Selection
We choose Qwen2.5-7B-Instruct (Team, 2024) as the base model, and first train it with SFT and then RL.
D.2 Contamination Detection Methods
Setup
For input question $q$, response $o$, and model $\pi_{\theta}$, we define the detection score as $f(q,o,\pi_{\theta})$. We treat the contaminated benchmark as the member set and the remaining uncontaminated half as the non-member set.AUROC is computed from the detection scores between the member and non-member set within a benchmark. We define the average log probability of model $\pi_{\theta}$ generating the response $o$ given the question $q$ as:
$\phi(q,o,\pi_{\theta})\;=\;\frac{1}{|o|}\log\pi_{\theta}\!\left(o\mid q\right).$
We provide experiments to illustrate why we choose the log probability on responses for most detection methods in appendix E.1, and assume that if models have seen the questions during the training, they would have more confidence during the generation, thus have higher log probabilities.
D.2.1 Generation-based Detection
Verbatim (Wu et al., 2025)
Verbatim-based approach (Wu et al., 2025) prompts the model to complete the remaining parts of a question based on partial prefixes.  (Wu et al., 2025) uses the partial-prompt completion rate measured by ROUGE-L (Lin, 2004), which calculates the overlap of the longest common subsequence between the generated and reference text. If the model memorizes the question during the training, the partial-prompt completion rate would be higher compared to unseen questions. We use 80% of the original problem to generate a partial completion. Using a lower ratio causes the LRM to answer the question directly instead of continuing the given sequence.
CDD (Dong et al., 2024)
CDD measures the variation of the generated responses, and assumes that if the responses share strong similarities, the question is more likely to be contaminated during the training.
$f(q,o,\pi_{\theta})=\sum_{d=0}^{\alpha\textperiodcentered l}\rho^{*}(d)=\sum_{d=0}^{\alpha\textperiodcentered l}\frac{\sum_{i=1}^{|G|}\mathbb{I}(\mathrm{ED}(o_{i},o_{\text{temperature=0}})=d)}{|G|}$
where ED represents Edit distance (Lcvenshtcin, 1966). We quantify the peakedness of the edit-distance distribution $\rho^{*}(d)$ by its cumulative mass within a similarity window, i.e., $F(d\leq\alpha l)$, where $\alpha\in[0,1]$ and $l=\max\{\operatorname{Len}(o)\mid o\in O\}$. In our experiments, we set $\alpha=0.5$, which performed better than other choices.
D.2.2 Perturbation-based Detection
Neighborhood (Mattern et al., 2023)
Neighborhood method calibrates the detection score $\phi(q,o,\pi_{\theta})$ with some unseen questions $q^{\prime}$ that share similar semantics to the original question $q$. We denote GPT-4o as $\mathcal{N}$ to augment the question $q$ and get $q^{\prime}\in\mathcal{N}(q)$. We use a total of five augmented samples to compute the detection score:
$f(q,o,\pi_{\theta})=\phi(q,o,\pi_{\theta})-\phi_{\mathrm{neighbor}}(q^{\prime},o,\pi_{\theta}),\qquad\phi_{\mathrm{neighbor}}(q^{\prime},o,\pi_{\theta})\;=\;\frac{1}{|\mathcal{N}(q)|}\sum_{q^{\prime}\in\mathcal{N}(q)}\phi(q^{\prime},o,\pi_{\theta}).$
D.2.3 Reference-based Detection
The detector assumes that they could access to the training distribution of the target model $\pi_{\theta}$, and have a reference model $\pi_{\text{ref}}$ to calibrate the detection scores. We choose bespokelabs/Bespoke-Stratos-7B (Labs, 2024) as the reference model $\pi_{\text{ref}}$ in all the experiments.
LiRA (Mireshghallah et al., 2022)
LiRA calibrates detection score by dividing.
$f(q,o,\pi_{\theta})=\frac{\phi(q,o,\pi_{\theta})}{\phi(q,o,\pi_{\text{ref}})}.$
Ref (Carlini et al., 2021)
Ref calibrates detection score by subtraction.
$f(q,o,\pi_{\theta})=\phi(q,o,\pi_{\theta})-\phi(q,o,\pi_{\text{ref}}).$
D.2.4 Reference-free Detection
LOSS (Carlini et al., 2021)
The detection score is as below:
$f(q,o,\pi_{\theta})=\phi(q,o,\pi_{\theta}).$
Zlib (Carlini et al., 2021)
Zlib calibrates the detection score with $\mathrm{Zlib}(o)$, which is a compression-based entropy/length proxy.
$f(q,o,\pi_{\theta})=\frac{\phi(q,o,\pi_{\theta})}{\mathrm{Zlib}(o)}.$
Min-K% (Shi et al., 2023)
Min-k% compute the detection score on k% tokens with lowest probabilities in the sequence. Following  (Shi et al., 2023), we choose k as 20 by default in all the experiments.
$f(q,o,\pi_{\theta})=\frac{1}{|\mathrm{min}\text{-}k(o)|}\sum_{i\in\mathrm{min}\text{-}k(o)}\Big[\log\pi_{\theta}\!\left(o_{i}\mid q,o_{<i}\right)\Big].$
Min-K++% (Zhang et al., 2024)
Min-k%++ standardizes the log probability before taking Min-K%. $\mu_{i}$ is the expectation of the next token’s log probbaility over the vocabulary $\mathcal{V}$ of the model $\pi_{\theta}$ given the prefix $q,o_{<i}$, and $\sigma_{i}$ is the standard deviation. We use k=20 by default.
$f(q,o,\pi_{\theta})=\frac{1}{|\mathrm{min}\text{-}k(o)|}\sum_{i\in\mathrm{min}\text{-}k(o)}\frac{\log\pi_{\theta}\!\left(o_{i}\mid q,o_{<i}\right)-\mu_{i}}{\sigma_{i}},$
$\mu_{i}=\mathbb{E}_{v\in\mathcal{V}}\!\left[\log\pi_{\theta}\!\left(v\mid q,o_{<i}\right)\right],\quad\sigma_{i}=\mathrm{Std}_{v\in\mathcal{V}}\!\left[\log\pi_{\theta}\!\left(v\mid q,o_{<i}\right)\right].$
Max-K% (Maini et al., 2024)
Max-k% compute the detection score on k% tokens with largest probabilities in the sequence. We use k=20 by default.
$f(q,o,\pi_{\theta})=\frac{1}{|\mathrm{max}\text{-}k(o)|}\sum_{i\in\mathrm{max}\text{-}k(o)}\Big[\log\pi_{\theta}\!\left(o_{i}\mid q,o_{<i}\right)\Big].$
D.3 Datasets
AIME 2024 & 2025
Olympiad-level mathematical-reasoning benchmarks consisting of 30 problems each from the American Invitational Mathematics Examination (AIME) 2024 and 2025.
AMC 2023
A high school level math benchmark consisting of 40 problems from the 2023 American Mathematics Competitions (AMC).
GPQA Diamond (Rein et al., 2024)
Graduate level scientific-reasoning, multiple-choice benchmark written by domain experts in biology, physics, and chemistry. The Diamond split is the hardest subset, with 198 questions retained only when expert annotators agreed, and non-expert baselines typically fail.
OlympiadBench (He et al., 2024)
An Olympiad level, bilingual, multimodal benchmark designed to test scientific reasoning in mathematics and physics. We use the English, text-only math subset consisting of 674 competition problems.
Minerva Math (Lewkowycz et al., 2022)
A challenging quantitative reasoning benchmark derived from Google’s Minerva work, consisting of 272 problems.
D.4 Implementation Details
SFT Implementation
We use the LLaMA-Factory (Zheng et al., 2024) implementation for our SFT experiments. By default, we adopt the SFT hyperparameters suggested by OpenThought3 (Guha et al., 2025) for medium dataset scales for the scenario that the base model evolves into LRMs, as shown below. To improve training efficiency, we employ FlashAttention-2 (Dao, 2023), DeepSpeed ZeRO-1 (Rasley et al., 2020), Liger kernels (Hsu et al., 2024), and asynchronous activation offloading (Daniel Han & team, 2023).
Training type Batch Size Context Length LR Epochs LR Scheduler Warmup Ratio Weight Decay Training Precision
Full 128 32,768 4e-5 5 cosine 0.1 0 bf16
For the extensive contamination to LRMs in the final stage, we use the hyperparameters as follows: Training type Batch Size Context Length LR Epochs LR Scheduler Warmup Ratio Weight Decay Training Precision Full 32 32,768 1e-5 7 cosine 0.1 0 bf16
RL Implementation
We use Verl (Sheng et al., 2024) implementation for our RL experiments. For RAFT and RAFT++, we follow the implementation of (Xiong et al., 2025). For GRPO, we follow DAPO (Yu et al., 2025) and do not introduce the KL term in our training in all the experiments. The detailed hyperparameters are as follows: Training type Batch Size Prompt Length Response Length $\epsilon$ LR Epochs Rollout Num Rollout Temp Training Precision Full 64 1,024 16,384 0.2 1e-6 1 4 0.6 bf16
Prompt template
We use a prompt template to enable long CoT during train on both SFT and RL. We use math template to AIME2024 & 2025, AMC2023, OlympiadBench (He et al., 2024), and Minervamath (Lewkowycz et al., 2022). We adapt multiple-choice template to GPQA Diamond (Rein et al., 2024).
Evaluation and Metric
We evaluate pass@1 and run 10 rollouts on AIME 2024 & 2025, AMC 2023, and 3 rollouts on OlympiadBench, GPQA Diamond, and Minerva Math to compute the pass@1. We use vLLM (Kwon et al., 2023) for the inference. All inference uses the same configurations: temperature=0.6, top_p=0.95, max_new_tokens=32,768.
[…]
1 was treated with pyridinium chlorochromate, forming product 2.
3 was treated with (dimethyl(oxo)-l6-sulfaneylidene)methane in DMSO at elevated temperature, forming product 3.
how many carbon atoms are there in product 3? A) 11, B) 10, C) 12, D) 14
Deduplication
We deduplicate our clean training datasets for both RL and SFT against the evaluation benchmarks using 13-gram overlap deduplication to ensure conclusive results.
Appendix E More Experiment Results
E.1 AUROC on Question, Question+Response, Thinking Process, and Non-thinking Process
We compare the AUROC computed on response tokens with computed on the question tokens, whole tokens, thinking tokens, and non-thinking tokens in the response. As shown in Tab. 7 and  10, none of the approach outperform the AUROC computed on the response tokens. Thus, we choose to compute all the detection scores on the response if applicable.
Table 7: AUROC (%) of contamination detection approaches using question tokens to compute the detection score, evaluated on the SFT-contaminated model w/o RL. $\Delta$ measures the difference with the reponse tokens as signal (Tab. 2). Results demonstrate that question tokens are not suitable for detection in the LRM contamination setting, compared with using response tokens.
Contamination Detection Methods Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Generation base
Verbatim (Wu et al., 2025) 47.58 49.86 47.56 53.56 52.52 65.50 52.76 +0.00
Perturbation base
Neighbor (Mattern et al., 2023) 47.04 56.10 44.22 32.44 49.99 54.62 47.40 -3.31
Reference base
LiRA (Mireshghallah et al., 2022) 49.78 43.12 42.22 46.00 48.77 55.00 47.48 -41.65
Ref (Carlini et al., 2021) 49.80 48.47 40.22 47.56 49.38 53.88 48.22 -17.28
Reference free
Zlib (Carlini et al., 2021) 48.15 50.78 60.00 40.89 54.23 52.75 51.13 -2.25
Min–K%++ (Zhang et al., 2024) 44.33 50.53 32.00 36.89 49.67 34.25 41.28 -8.33
Min–K% (Shi et al., 2023) 46.68 51.91 53.33 43.56 54.13 44.38 49.00 -15.96
Max–K% (Maini et al., 2024) 49.39 54.58 71.56 55.33 54.98 66.25 58.68 -11.15
Loss (Carlini et al., 2021) 47.38 54.03 58.44 46.00 55.61 52.00 52.24 -23.24
Table 8: AUROC (%) of contamination detection approaches using question + response tokens to compute the detection score, evaluated on the SFT-contaminated model w/o RL. $\Delta$ measures the difference with response tokens as signal (Tab. 2). Results demonstrate that considering both question and response tokens when computing detection scores actually harms the AUROC.
Contamination Detection Methods Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Perturbation base
Neighbor (Mattern et al., 2023) 54.56 43.42 48.89 38.44 56.09 61.50 50.48 -0.23
Reference base
LiRA (Mireshghallah et al., 2022) 79.31 72.49 89.56 71.11 66.07 89.00 77.92 -11.21
Ref (Carlini et al., 2021) 72.95 64.91 60.89 40.00 72.36 82.25 65.56 +0.06
Reference free
Zlib (Carlini et al., 2021) 46.96 54.24 69.33 41.11 43.31 42.00 49.49 -3.89
Min–K%++ (Zhang et al., 2024) 40.49 51.08 36.19 48.67 39.61 30.50 41.09 -8.52
Min–K% (Shi et al., 2023) 56.46 56.73 85.78 73.78 48.48 58.13 63.23 -1.73
Max–K% (Maini et al., 2024) 64.43 63.78 64.22 81.78 66.09 76.50 69.47 -0.36
Loss (Carlini et al., 2021) 59.18 58.44 87.11 76.44 49.75 63.12 65.67 -9.81
Table 9: AUROC (%) of contamination-detection methods using reasoning tokens only (tokens inside <think></think>) to compute the detection score, evaluated on the SFT-contaminated model w/o RL. The last column, $\Delta$, reports the difference relative to using the entire response as the signal (Tab. 2). Results show only minor differences between reasoning token only and whole-response signals, so we use the entire response in the main analysis.
Contamination Detection Methods Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Reference free
Zlib (Carlini et al., 2021) 49.67 59.60 70.89 35.24 48.60 46.00 51.67 -0.93
Min–K% (Shi et al., 2023) 70.76 66.46 82.89 74.52 70.34 82.25 74.54 +0.69
Max–K% (Maini et al., 2024) 67.40 62.00 66.67 82.38 65.87 79.00 70.55 +0.82
Loss (Carlini et al., 2021) 70.81 66.75 84.89 74.52 70.06 82.75 74.96 +1.21
Table 10: AUROC (%) of contamination-detection methods using non-reasoning tokens only (tokens after </think>) to compute the detection score, evaluated on the SFT-contaminated model w/o RL. The last column, $\Delta$, reports the difference relative to using the entire response as the signal (Tab. 2). Results show that using non-reasoning tokens degrades performance compared with whole-response signals, so we use the entire response in the main analysis.
Contamination Detection Methods Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Reference free
Zlib (Carlini et al., 2021) 59.73 51.49 49.11 61.67 63.37 74.25 59.94 +7.34
Min–K% (Shi et al., 2023) 59.26 54.82 48.22 80.00 64.8 76.5 63.93 -9.92
Max–K% (Maini et al., 2024) 60.86 54.93 55.33 69.52 59.76 77.5 62.98 -6.75
Loss (Carlini et al., 2021) 59.28 54.78 48.44 80.00 64.74 76.5 63.96 -9.79
E.2 Log-probability distribution before and after RL
We provide the log probability distribution for members vs. non-members in diverse scenarios. Fig. 56, and 7 demonstrate that further SFT and RAFT are unable to contract the distribution for members and non-members, while GRPO and RAFT++ could conceal the contamination evidence due to the PPO-style importance sampling/clipping term in the training objective.
Figure 5: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training on GPQA-Diamond. With additional GRPO or RAFT++ training on clean samples, the member and non-member log-probability distributions become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. In contrast, further RAFT training does not induce the earlier distribution collapse; as we explain in Sec. 3.2, the absence of a clipping term prevents it. Likewise, additional SFT does not collapse the membership distributions.
Figure 6: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training on OlympiadBench. With additional GRPO or RAFT++ training on clean samples, the member and non-member log-probability distributions become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. In contrast, further RAFT training does not induce the earlier distribution collapse; as we explain in Sec. 3.2, the absence of a clipping term prevents it. Likewise, additional SFT does not collapse the membership distributions.
Figure 7: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training on Minerva Math. With additional GRPO or RAFT++ training on clean samples, the member and non-member log-probability distributions become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. In contrast, further RAFT training does not induce the earlier distribution collapse; as we explain in Sec. 3.2, the absence of a clipping term prevents it. Likewise, additional SFT does not collapse the membership distributions.
E.3 AUROC for different training steps
We provide complete results of AUROC in different RL training steps. As shown in Tab. 11 and 12, we observe a monotonic decline in detection performance as the number of RL steps increases when using GRPO. While even with 156 steps in RAFT, there is no sign of AUROC decline. These results perfectly validate our theoretical analysis that RAFT is unable to conceal contamination, while GRPO could.
Table 11: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). The results demonstrate that additional GRPO steps better conceal the contamination evidence.
Contamination Detection Methods Steps Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Generation base
Verbatim (Wu et al., 2025) 64 45.46 50.49 48.89 57.56 52.74 59.63 52.46 -0.30
110 45.27 52.32 52.44 50.22 50.86 60.88 52.00 -0.76
156 46.51 51.09 53.33 53.56 52.84 58.50 52.64 -0.12
CDD (Dong et al., 2024) 64 55.47 51.08 43.33 60.00 60.18 62.00 55.34 -0.45
110 54.90 54.20 49.33 59.11 53.48 58.88 54.98 -0.81
156 53.99 43.42 34.22 70.89 56.32 62.38 53.54 -2.25
Perturbation base
Neighbor (Mattern et al., 2023) 64 54.10 39.68 50.67 44.22 53.42 60.5 50.43 -0.28
110 53.93 40.93 43.11 46.22 54.27 59.62 49.68 -1.03
156 53.32 41.07 47.11 38.67 54.49 60.75 49.24 -1.47
Reference base
LiRA (Mireshghallah et al., 2022) 64 74.41 84.65 70.22 87.78 81.04 82.75 80.14 -8.99
110 67.87 79.80 60.22 78.89 81.49 71.88 73.36 -15.77
156 66.41 77.28 54.22 80.22 77.62 75.13 71.81 -17.32
Ref (Carlini et al., 2021) 64 66.77 58.41 45.33 51.11 65.54 73.62 60.13 -5.37
110 64.92 54.70 44.44 52.22 66.47 72.37 59.19 -6.31
156 62.20 57.30 50.89 55.78 64.62 75.25 61.01 -4.49
Reference free
Zlib (Carlini et al., 2021) 64 45.94 54.99 66.22 35.56 46.65 39.38 48.12 -5.26
110 44.98 56.54 64.22 39.11 44.93 38.38 48.03 -5.35
156 44.82 53.21 63.11 41.33 46.43 39.12 48.00 -5.38
Min–K%++ (Zhang et al., 2024) 64 46.25 46.78 36.67 50.89 51.35 29.62 43.59 -6.02
110 43.17 44.69 28.33 50.00 49.44 46.25 43.65 -5.96
156 46.12 44.86 32.44 44.00 53.39 35.38 42.70 -6.91
Min–K% (Shi et al., 2023) 64 55.19 60.60 62.89 65.56 61.50 61.87 61.27 -13.69
110 49.98 58.49 61.78 61.33 56.20 49.00 56.13 -18.83
156 49.17 54.14 44.67 54.44 53.97 48.75 50.86 -24.10
Max–K% (Maini et al., 2024) 64 53.05 51.43 49.78 50.22 51.84 57.75 52.35 -17.48
110 51.02 51.39 49.78 53.33 51.88 47.50 50.82 -19.01
156 49.81 53.31 50.00 50.22 51.52 55.00 51.64 -18.19
Loss (Carlini et al., 2021) 64 55.22 60.50 62.44 65.78 61.50 62.12 61.26 -14.22
110 50.17 58.25 60.67 62.89 56.47 49.12 56.26 -19.22
156 49.32 54.04 44.22 54.22 54.20 48.50 50.75 -24.73
Table 12: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with RAFT in different steps. $\Delta$ measures the difference with the SFT contaminated model w/o RL (Tab. 2). Results demonstrate that even with more RL steps, RAFT is unable to conceal the contamination evidence.
Contamination Detection Methods Steps Olympiad GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
Generation base
Verbatim (Wu et al., 2025) 64 45.43 51.19 59.33 60.67 51.48 61.13 54.87 +2.11
110 45.67 51.24 58.67 58.67 52.33 57.50 54.01 +1.25
156 45.15 52.16 59.11 59.33 51.28 61.25 54.71 +1.95
CDD (Dong et al., 2024) 64 57.85 53.55 46.89 60.22 59.63 66.62 57.46 +1.67
110 55.97 52.86 43.11 52.89 56.45 58.25 53.26 -2.53
156 55.59 52.94 32.22 53.11 55.84 67.62 52.89 -2.90
Perturbation base
Neighbor (Mattern et al., 2023) 64 55.44 39.63 50.22 43.11 53.15 58.88 50.07 -0.64
110 55.06 38.47 46.89 40.67 54.20 63.38 49.78 -0.93
156 55.06 39.19 40.89 41.78 49.66 57.12 47.28 -3.43
Reference base
LiRA (Mireshghallah et al., 2022) 64 85.85 85.69 98.22 84.22 84.86 91.75 88.43 -0.70
110 84.32 85.76 94.44 88.44 84.55 91.25 88.13 -1.00
156 83.32 86.57 94.44 82.00 58.73 86.63 81.95 -7.18
Ref (Carlini et al., 2021) 64 76.02 65.94 74.67 46.00 73.99 75.62 68.71 +3.21
110 76.41 63.76 71.11 50.22 74.50 81.75 69.63 +4.13
156 75.92 64.23 67.33 54.67 56.73 81.38 66.71 +1.21
Reference free
Zlib (Carlini et al., 2021) 64 51.88 59.73 76.00 52.22 52.40 47.75 56.66 +3.28
110 52.39 62.20 76.22 49.78 74.51 45.88 56.29 +2.91
156 53.04 62.35 79.78 47.78 52.58 50.37 57.65 +4.27
Min–K%++ (Zhang et al., 2024) 64 51.14 54.13 51.78 62.00 59.86 52.75 55.28 +5.67
110 52.30 55.56 56.67 62.14 55.00 49.50 55.20 +5.59
156 51.26 59.72 61.19 58.44 52.53 63.12 57.71 +8.10
Min–K% (Shi et al., 2023) 64 72.28 69.56 88.44 86.67 71.63 78.88 77.91 +2.95
110 71.44 73.69 78.67 90.22 70.45 77.62 77.02 +2.06
156 70.76 71.12 84.00 80.67 56.12 79.38 73.68 -1.28
Max–K% (Maini et al., 2024) 64 68.11 68.37 68.44 90.67 69.21 80.50 74.22 +4.39
110 67.97 68.16 65.78 92.44 69.00 76.75 73.35 +3.52
156 69.14 68.30 69.33 90.67 58.06 76.00 71.92 +2.09
Loss (Carlini et al., 2021) 64 71.78 69.78 86.00 86.67 71.58 79.25 77.51 +2.03
110 71.09 73.54 78.44 91.11 70.13 77.12 76.90 +1.42
156 70.49 70.95 83.11 80.44 56.07 79.12 73.36 -2.12
E.4 Contamination Inflation Comparison
To demonstrate performance inflation from SFT contamination, we choose Qwen2.5-7B-Instruct as the base model and compare pass@1 across three training settings: (i) 15K clean samples from Openthought3 (Guha et al., 2025), (ii) 10K clean samples plus three full repetitions of six entire benchmark data (13,735 samples in total), and (iii) 1.2M clean samples (i.e., open-thoughts/OpenThoughts3-1.2M). As shown in Tab. 13, the contaminated model outperforms the 15K clean baseline by an average of 10.80% across six benchmarks and even surpasses OpenThoughts3-1.2M on GPQA-Diamond and Minerva Math datasets. These results demonstrate that benchmark contamination can easily yield substantial performance inflation.
Table 13: Pass@1 (%) of comparison between clean and contaminated SFT models. Bold=Best.
Training Data Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg.
15K clean 50.81 41.67 21.67 29.17 34.01 77.50 42.47
13.7K (clean ++ benchmarks) 58.52 59.09 40.00 40.00 44.49 77.50 53.27
1.2M clean 63.70 50.88 60.00 62.50 37.50 92.50 61.18
E.5 Log-prob distribution after extensive SFT contamination on LRMs
We provide the log prob distributions for members and non-members of deepseek distill models before and after contamination on Minvera Math and GPQA-Diamond in Fig. 8 and 9. Even though non-members have not been exposed to LRMs during the contamination, the log-prob would also increase, demonstrating that LRMs start to generalize after contamination.
(a) R1 Distill LLaMA results on Minerva Math
(b) R1 Distill Qwen results on Minerva Math
Figure 8: Log-prob distributions for members vs. non-members of advanced LRMs before and after contamination. After extensive SFT contamination on members, the log prob of both members and non-members increases at a similar margin, due to the generalization.
(a) R1 Distill LLaMA results on GPQA-Diamond
(b) R1 Distill Qwen results on GPQA-Diamond
Figure 9: Log-prob distributions for members vs. non-members of advanced LRMs before and after contamination. After extensive SFT contamination on members, the log prob of both members and non-members increases at a similar margin, due to the generalization.
E.6 Token Embedding Visualization of member and non-member of benchmarks
We provide visualization of token embeddings of member and non member of benchmarks. We can observe that we cannot distinguish between member and non-member using embeddings, which infer that embedding based method is hard to identify the membership. Specifically, we extract the last token from the last layer of the hidden space.
(a) Result on Minerva Math before SFT Contamination
(b) Result on Minerava Math after SFT Contamination
Figure 10: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
(a) Result on GPQA Diamond before SFT Contamination
(b) Result on GPQA Diamond after SFT Contamination
Figure 11: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
(a) Result on OlympiadBench before SFT Contamination
(b) Result on OlympiadBench after SFT Contamination
Figure 12: Last token embedding visualization of the final hidden layer before and after SFT contamination. In both cases, member and non-member embeddings are highly overlapped and therefore not distinguishable.
E.7 Additional Results
Table 14: Pass@1 (%) of the SFT-contaminated model after four additional epochs of SFT on clean data. The results show that further SFT does not make the model forget contamination; instead, pass@1 continues to increase by 0.25% across six benchmark on average compared to the SFT contaminated model, indicating persistent performance inflation.
Models Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg.
SFT Conta (w/o RL) + Further SFT 55.24 49.43 29.58 36.67 39.15 75.00 47.52
Table 15: Pass@1 (%) of advanced LRMs before and after SFT contamination with CoT on both clean and member data. Some LRMs with strong reasoning ability may not have a huge performance inflation after SFT contamination with CoT on both clean and members. Thus, we choose the SFT contamination with CoT on members only as the default setup in our main analysis.
Models Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg.
DeepSeek-R1-Distill-Llama-8B 52.10 43.94 33.33 43.33 32.97 84.58 48.38
$\hookrightarrow$ SFT w/ Clean & Mem 56.70 45.83 48.33 56.67 35.94 88.12 55.27
DeepSeek-R1-Distill-Qwen-7B 55.70 48.65 39.26 53.70 37.25 91.94 54.42
$\hookrightarrow$ SFT w/ Clean & Mem 55.81 45.08 40.00 55.42 37.87 88.12 53.72
Appendix F Computation resources
[…]